OpenStudy (anonymous):

What is the standard form of the quadratic function f(x) = -3(x -1)2 + 2?

OpenStudy (anonymous):

is it \[f(x)=-3(x-1)^{2}+2\] ??

OpenStudy (shaik0124):

f(x)=-3(x^2-1+2x)+2, =-3x^2+3-6x+2, = -3x^2-6x+6 is q eqn

OpenStudy (anonymous):

f(x)=-3(x^2-1+2x)+2, =-3x^2+3-6x+2, = -3x^2-6x+6 is q eqn